# CUE-RAG: Towards Accurate and Cost-Efficient Graph-Based RAG via Multi-Partite Graph and Query-Driven Iterative Retrieval

### v1

@@ 20      $\mathcal{E}_{e}\leftarrow\mathcal{E}_{e}\cup\{(v,k)\mid v\in\mathcal{V}_{k}\}$;
Algorithm 1 CUE-Index ($\mathcal{T},\alpha$)
@@ 6       $w_{i}\leftarrow\text{TokenLength}(t_{i})$;
3.3. Online Q-Iter
@@
In Algorithm 3, Q-Iter takes as input a query $q$ and a multi-partite graph index $\mathcal{G}$. The retrieval process operates under four parameters: the number of most relevant results $K$, maximum search depth $D$, beam size $M$ per depth level, and returned chunks size $N$. Q-Iter proceeds through three carefully designed stages as follows:
@@
Entity Anchoring. Inspired by cognitive theories of spreading activation (Anderson, 1983; Collins and Loftus, 1975), we first extract entities from the input query using LLM and identify their top-$K$ most similar entity nodes in $\mathcal{G}$ to form the initial query set $\mathcal{V}_{q}^{(0)}$. In parallel, we retrieve the top-$K$ knowledge units that are most semantically similar to the query and expand $\mathcal{V}_{q}^{(0)}$ by including all entities referenced in these units. This dual expansion incorporates both directly matched entities and entities introduced via semantically aligned knowledge units, enhancing the coverage and specificity of subsequent multihop retrieval. We refer to this combined process of Spreading Activation and Knowledge Anchoring as Entity-Anchoring, which yields the initial seed nodes $\mathcal{V}_{q}^{(0)}$. Due to space constraints, the full pseudo-code is deferred to Appendix A.
Iterative Retrieval. The algorithm 3 takes the seed nodes $\mathcal{V}_{q}^{(0)}$ as input. In Line 5, it initializes $\mathcal{Q}^{(0)}$ with tuples $(v,\bm{\phi}(q),\emptyset)$ for each anchor $v\in\mathcal{V}_{q}^{(0)}$, where $\bm{\phi}(q)$ is the query embedding and $\emptyset$ will accumulate the retrieved knowledge units. For each depth $d$ from $1$ to $D$, it processes the query queue $\mathcal{Q}^{(d-1)}$ in Lines 6-18. In Lines 8-9, for each $(v,\bm{\phi}_{v},\mathcal{S}_{v})\in\mathcal{Q}^{(d-1)}$, the algorithm retrieves top-$K$ new knowledge units $\mathcal{K}_{v}\subseteq\mathcal{V}_{\mathcal{K}}\setminus\mathcal{S}_{v}$ using $\bm{\phi}_{v}$, ensuring relevance while avoiding redundancy. In Lines 10-15, for each $k\in\mathcal{K}_{v}$, the algorithm first updates the query embedding by subtracting $k$’s embedding, dynamically shifting focus to uncovered information and avoiding redundancy (see Experiment 4.4). Next, the algorithm adds $k$ to $\mathcal{S}_{v}$, sorts it canonically, then uses re-ranker $\mathcal{R}$ to score $\oplus(\mathcal{S}_{v}^{\prime})$ for coherence with query $q$. Unlike semantic similarity measures, $\mathcal{R}$ specializes in query-context relevance scoring, allowing a more accurate assessment of knowledge units’ relevance to $q$. Then, the algorithm: (1) finds adjacent entities $\mathcal{V}^{\prime}$ via $k$’s edges as next-depth anchors, and (2) updates $\mathcal{C}$ with the new state tuple and its re-ranked $score$. Finally, in Lines 16-18, the algorithm prunes $\mathcal{C}$ to the top-$M$ tuples $\mathcal{C}^{\prime}$ based on their re-ranked $scores$, merges them into $\mathcal{C}^{*}$, and initializes $\mathcal{Q}^{(d)}$ for the next iteration.
Final Re-ranking. In Lines 19-23, the algorithm retrieves text nodes $\mathcal{T}^{\prime}$ citing $\mathcal{S}$ via $\mathcal{E}_{c}$, then selects top-$N$ by their re-ranked $scores$ to form final chunks $\mathcal{T}^{*}$ for augmented generation.
@@
2Input: query $q$, hierarchical graph index $\mathcal{G}=(\mathcal{V}_{\mathcal{T}}\cup\mathcal{V}_{\mathcal{K}}\cup\mathcal{V}_{e},\mathcal{E}_{c}\cup\mathcal{E}_{e})$, number of most relevant results $K$, search depth $D$, beam size $M$ per depth, returned candidate chunks size $N$
3Output: top-${N}$ relevant chunks $\mathcal{T}^{*}$
@@ 8       $\mathcal{C}\leftarrow\emptyset$;
11            $\mathcal{K}_{v}\leftarrow\mathop{\arg\max}\limits_{\begin{subarray}{c}k\in\mathcal{V}_{\mathcal{K}}\setminus\mathcal{S}_{v}\land(v,k)\in\mathcal{E}_{e}\land|\mathcal{K}_{v}|=K\end{subarray}}\cos\bigl{(}\bm{\phi}(k),\bm{\phi}_{v}\bigr{)}$
@@
13                   $\bm{\phi}^{\prime}_{v}\leftarrow\bm{\phi}_{v}-\bm{\phi}(k)$triangleright Update query embedding
14                  $\mathcal{S}_{v}^{\prime}\leftarrow\text{sort}(\mathcal{S}_{v}\cup\{k\})$triangleright Dictionary sorting
15                  $score\leftarrow\mathcal{R}(q,\oplus(\mathcal{S}^{\prime}_{v}))$triangleright Re-rank
16                  $\mathcal{V}^{\prime}\leftarrow\{v\mid(v,k)\in\mathcal{E}_{e}\}$
17                  $\mathcal{C}\leftarrow\mathcal{C}\cup\{(\mathcal{V}^{\prime},\bm{\phi}_{v}^{\prime},\mathcal{S}_{v}^{\prime},score)\}$
@@ 16                  $\mathcal{V}^{\prime}\leftarrow\{v\mid(v,k)\in\mathcal{E}_{e}\}$
20      $\mathcal{C}^{\prime}\leftarrow\mathop{\arg\max}\limits_{\begin{subarray}{c}(\mathcal{V},\bm{\phi},\mathcal{S},score)\in\mathcal{C}\land|\mathcal{C}^{\prime}|=M\end{subarray}}score$triangleright Pruning
21      $\mathcal{C^{*}}\leftarrow\mathcal{C^{*}}\cup\mathcal{C}^{\prime}$
22      $\mathcal{Q}^{(d)}\leftarrow\bigcup_{(\mathcal{V},\bm{\phi},\mathcal{S},score)\in\mathcal{C}^{\prime}}\{(v,\bm{\phi},\mathcal{S})\mid v\in\mathcal{V}\}$
@@ 21      $\mathcal{C^{*}}\leftarrow\mathcal{C^{*}}\cup\mathcal{C}^{\prime}$
27       $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{t\mid k\in\mathcal{S}\land(t,k)\in\mathcal{E}_{c}\}$
@@ 4. Experiments
In this section, we assess the effectiveness of CUE-RAG by addressing the following research questions:
  • Q1: How does CUE-RAG compare to existing baselines in terms of QA performance and cost efficiency?
  • Q2: How does our hybrid extraction strategy compare to alternatives and full LLM-based extraction in terms of QA performance and cost efficiency?
  • Q3: How does each stage of online retrieval Q-Iter contribute to the overall performance?
  • Q4: How sensitive is CUE-RAG to variations in hyperparameter settings?
All experiments are conducted on a Linux machine with an Intel Xeon 2.0 GHz CPU, 1024 GB of memory, and 8 NVIDIA GeForce RTX A5000 GPUs (each with 24 GB of memory), capable of running both LLaMA3.0-8B (Grattafiori et al., 2024) and Qwen2.5-32B (Bai et al., 2023) LLM models.
@@ 4.1. Experimental Setup
Datasets. We evaluate CUE-RAG’s capability primarily on three multihop QA benchmarks, MuSiQue (Trivedi et al., 2022b), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki. below) (Ho et al., 2020) dataset, all of which are extensively utilized within the QA and graph-based RAG research communities. Following prior work (Gutiérrez et al., 2024; Press et al., 2022; Trivedi et al., 2022a; Zhang et al., 2024), we extract 1,000 questions from each validation set, while the corresponding paragraphs for all sampled instances are preprocessed into a unified text corpora $\mathcal{T}$, maintaining question-paragraph alignment for evaluation. The details of these datasets are shown in Appendix Table 5.
Metrics. We primarily evaluate each solution’s performance along two dimensions: QA performance and cost efficiency. For QA performance, we report Accuracy (Acc. below) and F1 score (F1 below), where Acc. measures whether the golden answers are included in the generation answer rather than requiring exact match (Asai et al., 2023; Mallen et al., 2022), and F1 is computed via token-level overlap between golden and generated answers, balancing precision and recall to assess both answer completeness and correctness (Gutiérrez et al., 2024; Zhang et al., 2024). For cost efficiency, we quantify two metrics: total token expenditure for offline indexing and average token consumption per online query.
Baselines. We evaluate the performance of 13 solutions categorized as follows:
  • Simple Baselines: We consider two straightforward approaches: (1) the Zero-shot method, which directly applies an LLM for question answering without any retrieval, and (2) VanillaRAG, which retrieves relevant passages through query embedding similarity before using these passages as context for LLM answer generation.
  • Graph-Structured Baselines: We evaluate six graph-structured baselines that construct text-attribute KG using LLMs for content retrieval: HippoRAG (Gutiérrez et al., 2024), KETRAG (Huang et al., 2025), and local search of GraphRAG (LGraphRAG) (Edge et al., 2024) and three versions of LightRAG (LLightRAG for local search, GLightRAG for global search, and HLightRAG for hybrid search) (Guo et al., 2024), excluding GraphRAG’s global version as it targets abstract QA rather than specific multi-hop queries.
  • Tree-Structured Baselines: We evaluate two tree-based baselines: RAPTOR (Sarthi et al., 2024) (using semantic similarity) and SIRERAG (Zhang et al., 2024) (combining semantic and lexical similarity), both constructing multi-level trees for passage organization and retrieval.
  • Our proposed solutions: Our approach supports three operational modes, determined by a hyperparameter $\alpha\in[0,1]$, which controls the proportion of tokens in $\mathcal{T}$ processed by LLMs during offline indexing. CUE-RAG-0.0 ($\alpha=0.0$) demonstrates the strong performance of our LLM-free graph index and retrieval strategy. CUE-RAG-0.5 ($\alpha=0.5$) highlights the effectiveness of our hybrid extraction strategy, achieving strong performance while reducing 50% indexing cost. CUE-RAG-1.0 ($\alpha=1.0$) represents the full-capacity configuration, where all text chunks are processed by LLMs, enabling the system to achieve SOTA performance.
Settings. To ensure a fair comparison, all methods are implemented within a unified framework and environment (Zhou et al., 2025). During the LLM generation phase, we employ a strategy by duplicating the input query to the end of the original prompt. This design explicitly mitigates the ”lost in the middle” effect (Liu et al., 2023) observed in approaches like LightRAG, where placing the question before lengthy context can cause the model to overlook the task objective (Ram et al., 2023; Liu et al., 2023). Furthermore, inspired by the prompt designs of SIRERAG (Zhang et al., 2024) and KETRAG (Huang et al., 2025), which include phrases like ”Answer this question as short as possible” to prevent verbose or irrelevant outputs, we uniformly add this instruction to the final prompt of each method. This ensures concise responses, maintaining evaluation fairness and highlighting our method’s effectiveness. For other hyperparameters of each method, we follow the original settings in their available codes. For NLP pre-processing, we employ the following tools: cl100k_base from tiktoken library (OpenAI, [n. d.]) for word tokenization and counting, NLTK’s sent_tokenize (Bird et al., 2009) for sentence segmentation, and spaCy (Honnibal et al., 2020) for named entity recognition. For LLM inference, we use LLaMA3.0-8B (Grattafiori et al., 2024) and Qwen2.5-32B (Bai et al., 2023), with a maximum token limit of 8,000. In top-$K$ selection tasks, we set $K=5$ to accommodate token constraints. The embedding model for vector search is BGE-M3 (Chen et al., 2024), while re-ranking is performed using the lightweight yet powerful BGE-Reranker-v2-M3 (Chen et al., 2024). As for our proposed approach, CUE-RAG employs the following default configurations: the number of most relevant results $K$ is set to 3, search depth $D$ to 3, beam size $M$ per depth to 5, and returned candidate chunks $N$ to 5, ensuring that all methods ultimately retrieve the same number of query-relevant chunks (i.e., 5) for fair comparison. By default, we set the value of $\alpha$ to 1 if it is not explicitly specified. Additionally, BLEU score (Papineni et al., 2002) serves as the core selection metric for relevance scoring.
[…]
Figure 4. Token cost comparison during offline indexing using LLaMA3.0-8B. Reported costs include prompt and completion tokens for offline indexing. The seventh bar is empty because CUE-RAG-0.0 requires no LLM usage during indexing.
4.2. Performance Evaluation (Q1)
@@
Table 2. Overall Performance of RAG Solutions. The best and second-best results among the ten competitor solutions (excluding the CUE-RAG series) in each column are highlighted in bold and underlined, respectively.
@@
CUE-RAG-0.0 24.10 21.02 53.18 57.20 32.19 46.60 34.26 31.33 62.14 63.50 49.47 54.70 42.56 45.73
CUE-RAG-0.5 25.60 22.12 55.80 59.50 35.52 48.90 36.68 32.93 63.44 64.10 52.49 58.80 44.92 47.73
CUE-RAG-1.0 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
In the first set of experiments, we evaluate the QA performance and cost efficiency of CUE-RAG against ten existing solutions under three configurations: $\alpha$ equal to 0.0, 0.5, and 1.0, representing no LLM usage for indexing, selective LLM-aided indexing, and full LLM processing for indexing, respectively.
As shown in Table 2, CUE-RAG consistently achieves superior generation quality across multiple benchmark datasets when evaluated with different LLMs, with CUE-RAG-1.0 establishing SOTA performance in both average F1 and Acc. Specifically, it outperforms the strong baseline KETRAG by 21.53% in average F1, achieving up to a 73.15% improvement on the MuSiQue dataset when using Qwen2.5-32B, while also surpassing another competitive baseline SIRERAG by 5.75% in average Acc., with gains reaching up to 15.49% on the 2Wiki dataset when using LLaMA3.0-8B. Remarkably, CUE-RAG-0.5 demonstrates competitive QA performance and cost efficiency relative to KETRAG. It achieves improvements of 19.29%/27.19% in average F1/Acc., along with up to 74.00%/99.33% improvements in F1/Acc., respectively. Concurrently, it reduces token costs by up to 9.41%/83.87% during offline indexing and online querying shown in Figure 4 and 5. Notably, CUE-RAG-0.0, which entirely avoids LLM usage in indexing, still matches or exceeds the capabilities of SIRERAG and KETRAG. Compared to KETRAG, it improves average F1/Acc. by 13.01%/21.86% and F1/Acc. by up to 62.52%/89.65%, while reducing token costs by 100%/82.92% during offline indexing and online querying. Against SIRERAG, it achieves 52.50% higher average F1 and up to 101% improvement in F1, alongside 100%/56.06% reductions in indexing and querying costs.
Beyond outperforming the two strong baselines (KETRAG and SIRERAG), CUE-RAG demonstrates superior QA performance across all evaluated datasets when compared with other eight alternatives. This consistent advantage highlights the effectiveness of its innovative indexing and retrieval strategies in multi-hop QA scenery.
[…]
Figure 5. Token cost comparison per query during online retrieval with LLaMA3.0-8B. Reported costs include prompt and completion tokens for answer generation. For methods involving additional LLM-based query processing (e.g., keyword or entity extraction), such preprocessing costs are also included where applicable.
4.3. Effectiveness of Hybrid Extraction (Q2)

### v3

@@ 20 $\mathcal{E}_{e}\leftarrow\mathcal{E}_{e}\cup\{(v,k)\mid v\in\mathcal{V}_{k}\}$;
Algorithm 1 Clue-Index ($\mathcal{T},\alpha$)
@@ 6  $w_{i}\leftarrow\text{TokenLength}(t_{i})$;
3.2. Offline index: Clue-Index
We illustrate the two key stages of indexing as follows.
Stage-1: hybrid extraction. As shown in Algorithm 1, the hybrid extraction process begins by selecting a subset of core text chunks $\mathcal{T}_{c}\subseteq\mathcal{T}$ using Algorithm 2, based on a token constraint coefficient $\alpha\in[0,1]$. In Algorithm 2, we formulate the selection stage as a 0-1 knapsack problem, where each chunk’s relevance score determines its assigned value (line 5) and its token length corresponds to its weight (line 6), with the goal of selecting subsets to maximize total relevance under the token budget (line 9) — a classic optimization paradigm identical to 0-1 knapsack problem (Sahni, 1975).
In Algorithm 1, once $\mathcal{T}_{c}$ is identified (line 3), each text chunk $t_{i}$ is processed according to its selection status (lines 6–11). For $t_{i}\in\mathcal{T}_{c}$, an LLM extracts disambiguated knowledge units $\mathcal{K}_{i}$ that capture core semantics. For $t_{i}\in\mathcal{T}\setminus\mathcal{T}_{c}$, a lightweight NLP tool (Bird et al., 2009) is used to segment the text into sentences as knowledge units $\mathcal{K}_{i}$ to ensure broad coverage at lower computational cost. All text chunks and extracted knowledge units are instantiated as text nodes $\mathcal{V}_{\mathcal{T}}$ and knowledge unit nodes $\mathcal{V}_{\mathcal{K}}$, respectively, forming the upper and intermediate layers of the multi-partite graph. Finally, the Named Entity Recognition (NER) component in spaCy (Honnibal et al., 2020) is utilized to extract entities $\mathcal{V}_{k}$ for each knowledge unit $k$ (lines 12-13). All these extracted entities $\mathcal{V}_{k}$ are collected and form entity nodes $\mathcal{V}_{e}$, acting as anchors for precise information retrieval.
Stage-2: graph construction. As shown in Figure 3, Clue-Index is a multi-partite graph $\mathcal{G}$ built from a text corpora $\mathcal{T}$, organizing information across three semantic levels. In the upper layer, chunk nodes $\mathcal{V}_{\mathcal{T}}$ represent coarse-grained content by encoding text chunks. The intermediate layer contains knowledge unit nodes $\mathcal{V}_{\mathcal{K}}$, which capture semantically richer information distilled from the text. In the lower layer, entity nodes $\mathcal{V}_{e}$ encode fine-grained factual elements extracted from knowledge units. The graph is then connected by two types of edges: $\mathcal{E}_{c}$, linking each chunk node to its corresponding knowledge units, and $\mathcal{E}_{e}$, connecting knowledge units to the entities they mention. We present the construction algorithm in Algorithm 1.
3.3. Online retrieval algorithm: Q-Iter
@@
Given a query $q$ and a multi-partite graph index $\mathcal{G}$, our retrieval algorithm Q-Iter, outlined in Algorithm 3, identifies relevant information of $q$ from the corpora. It has four input parameters, i.e., the number of most relevant results $K$, maximum search depth $D$, beam size $M$ per depth level, and returned chunks size $N$. Q-Iter proceeds through three carefully designed steps as follows:
@@
Step-1: entity anchoring. The goal is to identify query-relevant seed nodes $\mathcal{V}_{q}^{(0)}$. Inspired by cognitive theories of spreading activation (Anderson, 1983; Collins and Loftus, 1975), we first extract entities from the input query using an LLM and identify their top-$K$ most similar entity nodes in $\mathcal{G}$ to form the initial query set $\mathcal{V}_{q}^{(0)}$. In parallel, we retrieve the top-$K$ knowledge units that are most semantically similar to the query and expand $\mathcal{V}_{q}^{(0)}$ by including all entities referenced in these units. This dual expansion incorporates both directly matched entities and entities introduced via semantically aligned knowledge units, enhancing the coverage and specificity of subsequent multihop retrieval. We refer to this combined process of Spreading Activation and Knowledge Anchoring as Entity Anchoring shown in Algorithm 4, which yields the initial seed nodes $\mathcal{V}_{q}^{(0)}$.
@@
2Input: $q$, $\mathcal{G}=(\mathcal{V}_{\mathcal{T}}\cup\mathcal{V}_{\mathcal{K}}\cup\mathcal{V}_{e},\mathcal{E}_{c}\cup\mathcal{E}_{e})$, $K$, $D$, $M$, NN
3Output: top-N{N} relevant text chunks $\mathcal{T}^{*}$
@@ 8  $\mathcal{C}\leftarrow\emptyset$;
11    $\mathcal{K}_{v}\leftarrow\mathop{\arg\max}\limits_{\begin{subarray}{c}k\in\mathcal{V}_{\mathcal{K}}\setminus\mathcal{S}_{v}\land(v,k)\in\mathcal{E}_{e}\land|\mathcal{K}_{v}|=K\end{subarray}}\cos\bigl{(}\bm{\phi}(k),\bm{\phi}_{v}\bigr{)}$;
@@
13        $\bm{\phi}^{\prime}_{v}\leftarrow\bm{\phi}_{v}-\bm{\phi}(k)$; $\triangleright$ Update query embedding
14       $\mathcal{S}_{v}^{\prime}\leftarrow\text{sort}(\mathcal{S}_{v}\cup\{k\})$; $\triangleright$ Dictionary sorting
15       $score\leftarrow\mathcal{R}(q,\oplus(\mathcal{S}^{\prime}_{v}))$; $\triangleright$ Re-rank
16       $\mathcal{V}^{\prime}\leftarrow\{v\mid(v,k)\in\mathcal{E}_{e}\}$;
17       $\mathcal{C}\leftarrow\mathcal{C}\cup\{(\mathcal{V}^{\prime},\bm{\phi}_{v}^{\prime},\mathcal{S}_{v}^{\prime},score)\}$;
@@ 16       $\mathcal{V}^{\prime}\leftarrow\{v\mid(v,k)\in\mathcal{E}_{e}\}$;
20 $\mathcal{C}^{\prime}\leftarrow\mathop{\arg\max}\limits_{\begin{subarray}{c}(\mathcal{V},\bm{\phi},\mathcal{S},score)\in\mathcal{C}\land|\mathcal{C}^{\prime}|=M\end{subarray}}score$; $\triangleright$ Pruning
21 $\mathcal{C^{*}}\leftarrow\mathcal{C^{*}}\cup\mathcal{C}^{\prime}$;
22 $\mathcal{Q}^{(d)}\leftarrow\bigcup_{(\mathcal{V},\bm{\phi},\mathcal{S},score)\in\mathcal{C}^{\prime}}\{(v,\bm{\phi},\mathcal{S})\mid v\in\mathcal{V}\}$;
@@ 21 $\mathcal{C^{*}}\leftarrow\mathcal{C^{*}}\cup\mathcal{C}^{\prime}$;
27  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{t\mid k\in\mathcal{S}\land(t,k)\in\mathcal{E}_{c}\}$;
@@
Figure 4. A toy example illustrating the online retrieval.
Step-2: iterative retrieval. The key idea is to progressively gather highly query-relevant and non-redundant knowledge units at each depth. Given the seed nodes $\mathcal{V}_{q}^{(0)}$, Algorithm 3 first initializes $\mathcal{Q}^{(0)}$ with tuples $(v,\bm{\phi}(q),\emptyset)$ for each anchor $v\in\mathcal{V}_{q}^{(0)}$ (line 5), where $\bm{\phi}(q)$ is the original query embedding that guides the semantic search and $\emptyset$ will accumulate the retrieved knowledge units. For each depth $d$ from $1$ to $D$, it processes the query queue $\mathcal{Q}^{(d-1)}$ (lines 6-18). For each $(v,\bm{\phi}_{v},\mathcal{S}_{v})\in\mathcal{Q}^{(d-1)}$, it retrieves top-$K$ new knowledge units $\mathcal{K}_{v}\subseteq\mathcal{V}_{\mathcal{K}}\setminus\mathcal{S}_{v}$ using $\bm{\phi}_{v}$, ensuring relevance while avoiding redundancy (lines 8-9). For each $k\in\mathcal{K}_{v}$, it first updates the query embedding by subtracting $k$’s embedding (lines 10-15), dynamically shifting focus to uncovered information and avoiding redundancy (This mechanism is termed Query Updating, as demonstrated in Experiment 4.4.). Next, the algorithm adds $k$ to $\mathcal{S}_{v}$, sorts it canonically, then uses re-ranker $\mathcal{R}$ to score $\oplus(\mathcal{S}_{v}^{\prime})$ for coherence with $q$. Unlike semantic similarity measures, $\mathcal{R}$ specializes in query-context relevance scoring, allowing a more accurate assessment of knowledge units’ relevance to $q$. Then, the algorithm: (1) finds adjacent entities $\mathcal{V}^{\prime}$ via $k$’s edges as next-depth anchors, and (2) updates $\mathcal{C}$ with the new state tuple and its re-ranked scorescore. Finally, the algorithm prunes $\mathcal{C}$ to the top-$M$ tuples $\mathcal{C}^{\prime}$ based on their re-ranked scoresscores, merges them into $\mathcal{C}^{*}$, and initializes $\mathcal{Q}^{(d)}$ for the next iteration (lines 16-18), where the re-ranker $\mathcal{R}$ guarantees $\mathcal{C}^{*}$ contain maximal evidential support for answering $q$ while maintaining computational efficiency.
Step-3: final re-ranking. The key idea is to retrieve the most query-relevant text chunks from $\mathcal{C}^{*}$ via re-ranking for augmented generation. The algorithm first retrieves text nodes $\mathcal{T}^{\prime}$ citing $\mathcal{S}$ via $\mathcal{E}_{c}$, and then selects top-$N$ by their re-ranked scoresscores to form final chunks $\mathcal{T}^{*}$ for augmented generation (lines 19-23).
Example 3.1.
Given a query $q=$ “When did the 1979-80 European Cup winner win the FA Cup?”, then Q-Iter executes the three steps:
Step-1 (entity anchoring): The LLM first extracts entities $\mathcal{V}_{q}^{(0)}$ = {1979-80 European Cup, FA Cup} from $q$. Simultaneously, it retrieves the most semantically similar knowledge unit “The European Cup was won by Nottingham Forest in the 1979-80 season” using the query embedding, expanding $\mathcal{V}_{q}^{(0)}$ to {1979-80 European Cup, FA Cup, Nottingham Forest}.
Step-2 (iterative retrieval): In the first iteration, the anchoring stage initializes with the seed nodes from Step 1. For each entity in $\mathcal{V}_{q}^{(0)}$, the algorithm retrieves the top-$K$ ($K=1$) semantically similar knowledge unit linked to it via the query embedding. For instance, FA Cup retrieves Chelsea F.C. won the FA Cup in 1970. After each iteration, Step 2 applies re-ranker $\mathcal{R}$ to score knowledge units against $q$. In this toy example, only the top-$M$ ($M=1$) relevant units are retained per iteration. The process terminates at maximum depth $D=3$. The complete retrieval path originates from Nottingham Forest to Forest won the FA Cup in 1898 and 1959.
Step-3 (final re-ranking): The algorithm maps the retrieved knowledge units back to their source text chunks, then re-ranks them based on $q$, and finally uses the most relevant chunks for RAG.
1
2Input: $q$, $K$, $\mathcal{G}=(\mathcal{V}_{\mathcal{T}}\cup\mathcal{V}_{\mathcal{K}}\cup\mathcal{V}_{e},\mathcal{E}_{c}\cup\mathcal{E}_{e})$
3Output: initial seed nodes $\mathcal{V}_{q}^{(0)}$
4# Spreading Activation
[…]
6foreach $v\in\mathcal{V}_{q}$ do
8 $\mathcal{V}_{q}^{(0)}\leftarrow\mathcal{V}_{q}^{(0)}\cup\mathcal{N}_{v}$;
9
10# Knowledge Anchoring
[…]
13return $\mathcal{V}_{q}^{(0)}$
Algorithm 4 Entity Anchoring($q,\mathcal{G},K$)
@@ 4. Experiments
In this section, we assess the effectiveness of Clue-RAG by addressing the following research questions:
  • Q1: How does Clue-RAG compare to existing baselines in terms of QA performance and cost efficiency?
  • Q2: How does our hybrid extraction strategy compare to alternatives and full LLM-based extraction?
  • Q3: What are the contributions of each step of Q-Iter?
  • Q4: How sensitive is Clue-RAG to hyperparameter settings?
@@ 4.1. Experimental Setup
Table 2. Overall performance of RAG solutions. The best and second-best results among the ten competitor solutions (excluding the Clue-RAG series) in each column are highlighted in bold and underlined, respectively.
@@
Clue-RAG-0.0 24.10 21.02 53.18 57.20 32.19 46.60 34.26 31.33 62.14 63.50 49.47 54.70 42.56 45.73
Clue-RAG-0.5 25.60 22.12 55.80 59.50 35.52 48.90 36.68 32.93 63.44 64.10 52.49 58.80 44.92 47.73
Clue-RAG-1.0 25.48 22.32 55.97 59.20 37.41 50.70 36.50 32.83 64.02 64.70 55.20 61.00 45.76 48.46
Datasets. We evaluate Clue-RAG on three multihop QA benchmarks: MuSiQue (Trivedi et al., 2022b), HotpotQA (Yang et al., 2018), and 2WikiMultiHopQA (2Wiki Below) (Ho et al., 2020). Following prior work (Gutiérrez et al., 2024; Press et al., 2022; Trivedi et al., 2022a; Zhang et al., 2024a), we use 1,000 validation questions per dataset, with corresponding paragraphs preprocessed into $\mathcal{T}$. See Appendix Table 5 for statistics.
Metrics. We evaluate each solution’s performance in terms of QA performance and cost efficiency. For QA performance, we report Accuracy (Acc. below) and F1 score (F1 below), where the former one measures whether the golden answers are included in the generation answer rather than requiring exact match (Asai et al., 2023; Mallen et al., 2022), and the latter one is computed via token-level overlap between golden and generated answers, balancing precision and recall to assess both answer completeness and correctness (Gutiérrez et al., 2024; Zhang et al., 2024a). For cost efficiency, we quantify two metrics: total token expenditure for offline indexing and average token consumption per online query $q$.
Baselines. We test the following 13 solutions in 4 categories.
  • Simple baselines: We consider 2 straightforward approaches: (1) the Zero-shot method, which directly applies an LLM for QA without any retrieval, and (2) VanillaRAG, which retrieves relevant passages through query embedding similarity before using these passages as context for generation.
  • Graph-based baselines: We evaluate 6 graph-based baselines that construct text-attribute KG using LLMs for content retrieval: HippoRAG (Gutiérrez et al., 2024), KETRAG (Huang et al., 2025), and local search of GraphRAG (LGraphRAG) (Edge et al., 2024), and 3 versions of LightRAG (LLightRAG for local search, GLightRAG for global search, and HLightRAG for hybrid search) (Guo et al., 2024), excluding GraphRAG’s global version as it targets abstract QA rather than multi-hop QA.
  • Tree-based baselines: We evaluate 2 tree-based baselines: RAPTOR (Sarthi et al., 2024) and SIRERAG (Zhang et al., 2024a), both constructing multi-level trees for passage organization and retrieval.
  • Our proposed solutions: Our approach supports three operational modes, determined by a hyperparameter $\alpha\in[0,1]$, which controls the proportion of tokens in $\mathcal{T}$ processed by LLMs during offline indexing. Clue-RAG-0.0 ($\alpha=0.0$) demonstrates the strong performance of our LLM-free graph index and retrieval strategy. Clue-RAG-0.5 ($\alpha=0.5$) highlights the effectiveness of our hybrid extraction strategy, achieving strong performance while reducing 50% indexing cost. Clue-RAG-1.0 ($\alpha=1.0$) represents the full-capacity configuration, where all text chunks are processed by LLMs, enabling the system to achieve SOTA performance.
Settings. All experiments are conducted on a Linux machine with an Intel Xeon 2.0 GHz CPU, 1024 GB of memory, and 8 NVIDIA GeForce RTX A5000 GPUs (each with 24 GB of memory), capable of running both LLaMA3.0-8B (Grattafiori et al., 2024) and Qwen2.5-32B (Bai et al., 2023) LLM models. To ensure a fair comparison, all methods are implemented under a unified framework (Zhou et al., 2025) using their default configurations. Additional experimental details are provided in Appendix A.1.
4.2. Overall performance evaluation (Q1)